Shift report 20030307 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Mar 07 2003 - 15:56:50 EST

  • Next message: Apache: "Shift report 20030307 16:00-24:00"
    ------------------------------------------------------------
    Date: Fri Mar  7 15:56:50 2003, Shift: 08:00-16:00
    Supervisor: ykl, with: 
    Log book#14 pages 15 - 16
    
    Shift Summary:
    ------------------------------------------------------------
    1.The beams is off since 9.30AM. They are beginning to inject now.
    
    2.MRS @90 deg, A1050, and FS @4 deg 2/3A.
      run 8246 had 57k trig3 and 11.7K trig 2.
      run 8247(DVM), run 8248(pedestals), run 8249(TPC pedestals).
    
    3.D5 tripped. Fan for the shunt was replaced.
      T3-C3-yv tripped and is now off.
    
    Plan: MRS @ 90 deg, A1050, for lots of statistics;
          FS @ 4 deg, 1/1A for 100K trig 2.
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Fri Mar 07 2003 - 16:01:25 EST